Starting from the same point at the same time Samuel and Heather rode their mini-bikes in the same direction. Samuel traveled 3 Km/h faster than Heather. After 3.5 hours Samuel stopped for a half-hour break, but Heather kept going. At the end of Samuel’s break, Heather was 8 km ahead of him. How fast was each going?
Answer by ankor@dixie-net.com(22740)

Starting from the same point at the same time Samuel and Heather rode their mini-bikes in the same direction.
Samuel traveled 3 Km/h faster than Heather.
After 3.5 hours Samuel stopped for a half-hour break, but Heather kept going.
At the end of Samuel’s break, Heather was 8 km ahead of him.
How fast was each going?
:
let s = Heather's speed
then
(s+3) = Samual's speed
:
From the information given we know S traveled for 3.5 hr and H traveled for 4 hrs
:
Write a distance equation; dist = time * speed
H's dist - S's dist = 8 km
4s - 3.5(s+3) = 8
4s - 3.5s - 10.5 = 8
.5s = 8 + 10.5
.5s = 18.5
Multiply both sides by 2
s = 37 km/hr is H's speed
and obviously
39 + 3 = 40 km/hr is S's speed
;
:
Check this find the actual dist each traveled
4 * 37 = 148 km
3.5*40 = 140 km
--------------
differ: 8 km